Question #7: Modes of inhibition determine the modes of inhibition of two inhibitors on an enzyme by using excel to graph the effects of the different inhibitors on the enzyme.

[S] (mM)                        No inhibitor                  2 mM Inhibitor #1            100 mM Inhibitor #2

3

10.4

4.1

2.1

5

14.5

6.4

2.9

10

22.5

11.3

4.5

30

33.8

22.6

6.8

90

40.5

33.8

8.1

  1. What are the values of Vmax and Km in the absence of each inhibitor and in the presence of both inhibitors?
  2. What type of inhibition does each inhibitor exhibit?
Answer #1
[S] mM 1/[S] 1/mM V without inhibitor mM/min 1/v without inhibitor min/mM V with inhibitor 1 1/V with inhibitor 1 V with inhibitor 2 1/V with inhibitor 2
3 0.33 10.4 0.09615 4.1 0.2439 2.1 0.4762
5 0.2 14.5 0.06897 6.4 0.1563 2.9 0.3448
10 0.1 22.5 0.04444 11.3 0.0885 4.5 0.2222
30 0.033 33.8 0.02959 22.6 0.0448 6.8 0.1471
90 0.011 40.5 0.02469 33.8 0.0296 8.1 0.1235

The equation of above Double reciprocal plot without inhibitor is :

y = 0.226x+0.022 (y = mx+c)

This can be compare to an equation 1/Vo = (Km/Vmax)1/[S] + 1/Vmax

Hence slope, m = Km/Vmax = 0.226

c = Y-intercept = 1/Vmax = 0.022

Vmax = 1/0.022 = 45.45 mM/min

Slope, m = Km/Vmax = 0.226

Km = 0.226 x 45.45 = 10.4977 mM

The equation of above Double reciprocal plot with inhibitor 1 is :

y = 0.671x+0.022 (y = mx+c)

This can be compare to an equation 1/Vo = (Km/Vmax)1/[S] + 1/Vmax

Hence slope, m = Km/Vmax = 0.671

c = Y-intercept = 1/Vmax = 0.022

Vmax = 1/0.022 = 45.45 mM/min ( It remained same)

Slope, m = Km/Vmax = 0.671

Km = 0.671 x 45.45 = 30.497 mM (It was increased)

The type of inhibition with inhibitor 1 is Competitive inhibition because Vmax is unaffected and Km is increased.

The equation of above Double reciprocal plot with inhibitor 1 is :

y = 1.118 x+0.112 (y = mx+c)

This can be compare to an equation 1/Vo = (Km/Vmax)1/[S] + 1/Vmax

Hence slope, m = Km/Vmax = 1.118

c = Y-intercept = 1/Vmax = 0.112

Vmax = 1/0.112 = 8.929 mM/min ( It was decreased)

Slope, m = Km/Vmax = 1.118

Km = 1.118 x 8.929 = 9.982 mM (It was decreased)

The type of inhibition with inhibitor 1 is Uncompetitive inhibition because both Vmax and Km are decreased.